please help me :find the improper fraction with a denominator of 6 that is equivalent to 5 1/2

Answer by nyc_function(2741)

x/6 = 5 1/2
x/6 = 11/2
2x = 66
x = 66/2
x = 33
The fraction you need is 33/6, which is = 5_1/2.
ANSWER: 33/6